Prove the following sentences using any prove method [ the most appropriate one]:

a) If r is rational and s is irrational, prove that 2r+s is irrational.

b) If t and s are integers and t × s is even, then t is even or s is even.


1
Expert's answer
2021-08-19T15:15:09-0400

Solution:

(a):

Suppose r is rational and s is irrational. Assume for contradiction that r + 2s is rational. Then there are integers a and b with b "\ne" 0 such that r + 2s = "\frac ab" . Since r is rational, there are integers c and d such that r ="\frac cd" . Then 2s = "\frac ab-\frac cd" = "\frac{ad−bc}{bd}" is rational, contradicting the fact that 2s is irrational.

Hence, proved.

(b):

Given that t and s are integers and t × s is even.

So, t × s = 2m, for some constant m.

"\dfrac{t × s }2=m
\\\Rightarrow \dfrac{t }2× s=m \ or\ \dfrac{s }2× t=m"

"\Rightarrow" Either t is divisible by 2, so it is even

or s is divisible by 2, so it is even.

Hence, proved.